OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve 1x^2+4xy+4y^3=41 at the point (1,2)

OpenStudy (danjs):

hi

OpenStudy (anonymous):

hello! I am not sure how to do this problem

OpenStudy (danjs):

ok one sec

OpenStudy (danjs):

the first term is simply a power rule wrt x

OpenStudy (danjs):

the second term, is using the product rule and the chain rule \[4x \frac{ d }{ dy }(y)\frac{ dy }{ dx } + \frac{ d }{ dy }(4y ^{2})\frac{ dy }{ dx }\]

OpenStudy (danjs):

the third term is a chain rule: \[\frac{ d }{ dy }(3y ^{2})\frac{ dy }{ dx }\]

OpenStudy (danjs):

You get all that part?

OpenStudy (anonymous):

yes I get it so far

OpenStudy (danjs):

now you have to solve that equation for dy/dx

OpenStudy (danjs):

let dy/dx = y'

OpenStudy (danjs):

2x + 4x*y' + 12y^2 *y' = 0

OpenStudy (danjs):

solve that for y'

OpenStudy (danjs):

then use the point they gave you to find the slope dy/dx at that point

OpenStudy (anonymous):

ok thank you i am going to try it